OpenStudy (anonymous):

Find the value of x + y if 3x - ix + 2y + yi = 5 + 4i, where i = √-1

OpenStudy (anonymous):

@apoorvk Please help?

OpenStudy (experimentx):

equate the real values and imaginary values of both sides

OpenStudy (apoorvk):

Okay, just compare the real an imaginary parts: 3x + 2 y = 5 .....{real} and (-x + y)i = 4i or, -x + y = 4

OpenStudy (apoorvk):

solve these two linear equations, and find out 'x' and 'y'!

OpenStudy (anonymous):

Oh ok! I think I go tit. 3x + 2y = 5 2x - 2y = -8 x = -3/5 y = 3.4 3.4 - 0.6 = 2.8?

OpenStudy (apoorvk):

Yup, right-O!!!

OpenStudy (anonymous):

Thanks! I just didn't know that we could do that.